OpenStudy (anonymous):

Which of the following system has no solution: A) 3x -2y =1 6x -4y= 6 B) -x+ y = 1 2x - 3y = 4 C) 4x + y = 7 x + 4y = 4 D) 3x- 4y = 2 -2x+ 4y = 9

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

A) 3x -2y =1 6x -4y= 6

OpenStudy (anonymous):

here a1/a2 = b1/b2 <>c1/c2

OpenStudy (anonymous):